If logb 3 = 0.8 and logb 6 = 1.3, find the value of logb 18
Answer by stanbon(75887)

If logb 3 = 0.8 and logb 6 = 1.3, find the value of logb 18
-----------------
logb 18 = logb (3*6) = logb 3 + logb 6 = 0.8 + 1.3 = 2.1
---------------
Cheers,
Stan H.